OpenStudy (anonymous):

John was visiting three cities that lie on a coordinate grid at (-4, 5), (4, 5), and (-3, -4). If he visited all the cities and ended up where he started, what is the distance he traveled? Round your answer to the nearest tenth. (like 3.2 or 5.7)

OpenStudy (anonymous):

Find the distance between the points using the distance formula \[\large \bf d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would it be 16 ?

OpenStudy (solomonzelman):

the points (-4, 5) and (4, 5) lay on the same line y=5, so the distance between these points is just: 4-(-4) So you know the first side of 8 units. You do however need the distance formula for two other sides. 1. Distance between (-3, -4) and (-4, 5) 2. Distance between (-3, -4) and (4, 5)

OpenStudy (solomonzelman):

Distance between (-3, -4) and (-4, 5) = \(\small\color{black}{ \displaystyle \sqrt{(-4-5)^2+(-3--4)^2}=?}\) ------------------------------------------------------- Distance between (-3, -4) and (4, 5) = \(\small\color{black}{ \displaystyle \sqrt{(-4-5)^2+(-3-4)^2}=?}\)

OpenStudy (solomonzelman):

after you find all sides, add them up

OpenStudy (anonymous):

would it be 16?

OpenStudy (anonymous):

I had gotten 12 but I know definitely it isn't the answer for the question